slope of x/-2 + y/8 = 1
Answer by rapaljer(4671)

I recommend that you solve for y in order to write the equation in slope-intercept form.
x/-2 + y/8 =1
Multiply both sides by 8 to clear the fractions:
-4x + y = 8

Add +4x to each side:
y=4x+8

y=mx+b

The coefficient of x in this form is the slope. Slope m=4.

For additional explanation on Slope of a Line, please see my own website. To find my website, click on my tutor name "rapaljer" anywhere in algebra.com. Click on the link there that will take you to the Homepage of my website, which is a subwebsite at Seminole State College of Florida. From my "Homepage", look for the link "Basic, Intermediate, and College Alegebra: One Step at a Time." Click on "Basic Algebra", and look in "Chapter 4" for several topics on Graphing and Slope. Here you will find my own complete non-traditional, user-friendly explanation of the topics that my own students found MUCH easier to understand than the traditional textbooks! Many of my exercises here are explained in great detail in the MATH IN LIVING COLOR pages that go with this curriculum. Best of all, it's all FREE!
Check also for my VIDEO on "Graphing and Slope of a Line" that was made a few years before I retired of me actually teaching this topic. To see the videos, from my Homepage, look for "Rapalje Videos in Living Color." Then click on "Basic Algebra" and look for the topic on mentioned above. If you have RealPlayer installed on your computer, then this is a FREE download. If you do not have RealPlayer, then you will have to download RealPlayer, but this too is a FREE download.

I have explanations on nearly every topic in algebra and many from geometry as well, and many videos from my younger days as well. Everything is FREE, and I don't even have advertising on the pages!! If you don't need help with math, then visit my website to see the JOKE PAGE--all clean jokes, safe for the little ears in your family!!!

Dr, Robert J. Rapalje, Retired
Seminole State College of Florida